OpenStudy (anonymous):

find an equation of the line that satisfies the given condition. Through (8,7) and (7,9)

OpenStudy (luigi0210):

First find the slope: \[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] Then use the point slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] Then just get it into slope-intercept form \[\LARGE y=mx+b\]